# An Exploration of Default Images in Text-to-Image Generation

### v4

@@ 5.1. Default Images
Figure 4 showcases the ten default images identified in our manual study, based on the set of input prompts from Section 4.1. Some default images occurred more frequently, as depicted in Figure 6. Figure 5 showcases how the same default image (in this case, ‘Lady-Birdhead’) appears in response to different prompts, with only minor visual variations. Due to Midjourney’s internal diversity sampling mechanism, the relationship between prompt and default image is many-to-many: one prompt can produce multiple different default images, and the same default image can appear in response to multiple different prompts (cf. Figure 2 and Figure 7B).
Figure 5. An example of how dissimilar prompts (listed below each image) generate the same default image in Midjourney v6 and v6.1, with only slight variations.
[…]
@@
Figure 6. Frequency of occurrence of default images in our experiments. A1–A6 denote the different test sets, comprising 130 prompts (Section 4.1 and Appendix A).
[…]
@@
The default images we identified exhibit recurring motifs, reflecting the model’s aesthetic biases when faced with ambiguous or unrecognized prompts. Human figures and natural elements – such as birds, clouds, trees, and flowers – appear as primary motifs, separately or merged into strange creatures. Human-vegetal and/or human-animal fusions emerge as key subjects in default images, appearing across different atmospheres, from oneiric scenes to mythological metamorphoses. Some identified default images confirm that similarly to human dreams, “AI has an uncanny ability to combine seemingly disparate elements into sometimes sinisterly coherent representations” (Martín Prada, 2025).
@@
Although the literature on AI imagery attests that visual expression tends to vary according to specific TTI generators, and some note ‘default styles’ associated with different models (Manovich, 2024), the curated series of 10 images presented in Figure 4 reveals significant aesthetic disparities. While the five images in the lower row are reminiscent of abstract expressionist and surrealist painting, also evoking magical realism and epic fantasy films, the first five images draw from other artistic movements. As shown in figures 4 and 5, the female portrait is also an important theme, particularly prone to formal and stylistic derivations. From the solitary woman standing before a romantic landscape – in which Caspar David Friedrich’s painting resonates – to close-up representations that seem to combine the tradition of portrait painting with lyrical abstraction and contemporary fashion photography, these images illustrate the hybrid and composite nature of AI-generated pictures. Figure 5 further emphasizes how AI-generated pictures can condense multiple artistic references into a single visual composition that, similarly to a musical piece, can then unfold in subtle variations. This series reflects the wide dissemination of individual portraits, both through the digitization and online circulation of art collections – including popular museum masterpieces – and through the massive production of born-digital pictures that fuel social media.
@@
The intriguing association of a stereotyped female figure and a bird in place of a hat poses a twofold question concerning the generation of default images. Apparently, the TTI model was unable to understand the prompt or correctly recognize some of the images it was trained with. Feather hats are a recurring theme in the history of female portraiture and fashion since the Renaissance, from Titian’s Portrait of a Young Woman (c. 1536) to Rubens’ Portrait of Helena Fourment (c. 1630), from Klimt’s The Black Feathered Hat (1910) to Karl Lagerfeld’s exuberant designs for Chanel’s Spring Collection in 1992. The examples are countless, hence it seems plausible that many of these images were used to train the model. However, there seems to be a fundamental difference between human perception of visual representations of feather hats and the way computer vision ‘sees’ such images. Rather than a dataset bias, what is at stake here is a perceptual bias (Offert and Bell, 2021). Following a metonymical logic (i.e., taking a part to stand in for the whole), the model takes the part for the whole, assuming that the feathers correspond to the presence of a bird whose species differs across the variations of the default image. This is an insightful demonstration of “the nomadic existence to which images seem to be condemned within the historical formations of knowledge” (Alloa, 2021).
@@ 5.2.1. Effects of seed values
We visually analyzed how different seed values affect the default images. Using different seeds with the same prompt resulted in different default images, partially sharing some of the motifs (Figure 7–A1). While there are similarities, one might not associate these images with each other. Thus, the best way to generate default images is by having a fixed seed value for all image generation runs. The images in Figure 7–A2 show common default image motifs, including portraits of women and animals (mostly birds).
Figure 7. Ablation studies: (A) Default images generated with different seed values (A1) and image similarities with variations in both seed values and prompts (A2). (B) Default images created with different style modifiers (seed 123456 and model version 6.1). (C) The examples at the top demonstrate how known terms (computer, browsing) overpower the unknown terms (in bold), producing images that are not default images. The two images at the bottom demonstrate how motifs from default images can be created in specific contexts (e.g., on a table). (D) Default images generated using the prompt “hevonen” and seed 123456 with Midjourney model version 6.1, 6.0, 5.2, and 5.1. Similarities in default images point to similarities in model training data. The four different images generated in response to a single prompt are the result of Midjourney’s latent space sampling and internal image diversity mechanisms.
[…]
@@ 5.2.2. Effects of style modifiers on default images
We found default images respond to style modifiers just like regular images. The same default image motifs reliably arise with different styles. Figure 7–B depicts how the prompt “kinkku”, which triggers default images, can be modified with different styles, like “in the style of cubism” or “oil on canvas”.
@@ 5.2.4. Effects of combining multiple default image prompts
We found that when we combined two of the default image prompts (from either the same or different sets), they would still reliably yield default images. As the number of prompts increased, the number of resulting default images appeared to decrease. When we combined all 20 prompts from a set, the default images identified in the main study (see Figure 4) were rare.
@@ 5.2.5. Effects of model versions
Last, we analyzed how different model versions affect the default images. In our experiments with Midjourney, both model versions 6.0 and 6.1 consistently generated very similar default images with the same prompt, as depicted in Figure 7–D. Older model versions (e.g., 5.1 and 5.2) generated default images with a more surreal aesthetic, but with similar motifs, including floating objects, human portraits intertwined with natural elements, either plants or animals, or both.
Figure 8. Responses from the online user study ($N=48$) demonstrate how default images can lead to user dissatisfaction.
[…]
@@ 5.3.2. Results
Figure 8 depicts the results of our user study, showing that unknown terms, which potentially trigger default images, negatively affect user satisfaction, but not in all cases.
@@ 5.4. Default Images on Midjourney
Figure 9. Selected examples of default images and prompts identified on Midjourney.
[…]
@@ 5.4.1. Results
Our computational analysis of images on Midjourney resulted in the identification of 4,710 clusters of potential default images. In total, the clusters comprise 35,913 images, representing 4.8% of our dataset. This percentage refers only to our filtered dataset and parameters (i.e., less than 25 characters per prompt, lexical diversity $\geq$ 0.5, and semantic similarity << 0.3) and should not be interpreted as representative of the overall prevalence of default images on Midjourney. The mean number of images per cluster is 7.62 (SD = 8.72, Median = 5, Min = 3, Max = 96). The lexical diversity of prompts within clusters ranges from 0.5 to 1.0 (Mean = 0.92, SD =0.12, Median = 1.0), and the semantic similarity from 0.01 to 0.3 (Mean = 0.2, SD =0.05, Median = 0.2). From smaller-scale experiments, we note the largest clusters were stable across multiple parameter settings. Some selected canonical examples are depicted in Figure 9 and Appendix C.
@@
  • We found evidence of short prompts like the ones in our manual study being in real-world use by Midjourney users. These prompts include, for instance, fantasy names, domain names, oxymorons, conceptual metaphors, abstract noun phrasing, poetic paradoxes and lyrical abstraction, surreal compound phrases, and other neologisms and creative compounds. This confirms that the six categories of prompts likely triggering default images in our manual study are realistic and have practical relevance.
  • The clusters contain recurrent themes identified in our earlier manual study, such as portrait-style images, intricate cityscapes, and dream-like imagery. This includes some of the iconic Midjourney motifs, such as a person walking up an eternal staircase or road, a town on a floating island, abstract and diffuse color arrangements, and depictions of seemingly the same person.
  • Major versions of Midjourney have different sets of default images (as expected), although some versions (such as version 3 and 4) share common default image motifs. This is a result of training data being shared between training runs.
@@ 6.1. Default Images in Text-to-Image Generation
Default images in TTI generation are an underexplored phenomenon. We presented the first study on default images in Midjourney. When prompts are vague and contain unknown terms, Midjourney may generate a default image, suggesting that the model defaults to familiar visual elements when it encounters words outside of its training data.
Our computational analysis of images on Midjourney reveals that default images are not a theoretical construct, they occur daily in response to real-world creative user prompts. Some of the images we identified in our analysis represent iconic Midjourney failure modes that many users have encountered, especially with the early Midjourney versions. Our findings confirm that these default images are not generated randomly and, on the basis of our experiments, have a relatively narrow thematic and formal range. Uncanny elements emerged in common motifs that often depicted human portraits (usually women) and animals (especially birds), along with other botanical or organic forms. Earlier versions of Midjourney have a different set of default images, often including motifs such as balloons, butterflies, fishes, floating islands, towns, and intricate compositions (see Figure 1 and Appendix C). Although many default images appear to share a common style, this straightforward association may obfuscate a more complex reality. For instance, in model version 5, some default images appear related to surrealism and cartoonish grotesque (cf. Figure 1), with other identifiable references from art history, namely Dutch and Flemish Renaissance painting, particularly Hieronymus Bosch and Bruegel the Elder. AI-generated images are hybrid and composite, condensing multiple forms and references learned from the training dataset into a single visual output. As Wasielewski points out “style is a gradient”; therefore, an AI image is unlikely to correspond to a specific artistic movement, rather reflecting “a fluctuating set of visual characteristics with highly fuzzy boundaries” (Wasielewski, 2023).
Our findings also demonstrate that default images may negatively affect user satisfaction. In the following section, we describe the characteristics of default images.
@@ 6.3. Explainability and Alignment of Text-to-Image Generation Models
What does ‘absolute ownership’, ‘nallopand’, or a ‘protruding goofus’ look like? Humans could perhaps imagine some creative imagery for these abstract concepts and lexical inventions. However, machines struggle to produce a result (cf. Figure 1). This is not surprising, given that these terms were likely not part of the TTI models’ training data. While polysemia and ambiguity will likely continue to be an issue in the alignment of TTI models with human intent (consider, for instance, the polysemous prompt “a plant on a bank”), default images are a different phenomenon – a type of image that appears in response to prompts that are not understood by the model, revealing a misalignment between model and human intent.
@@
Daras and Dimakis experimented with how gibberish prompts are understood by the model and suspected that TTI models have developed a ‘secret language’ (Daras and Dimakis, 2022). Their hypothesis was that DALL-E 2 had made ‘hidden’ associations between certain terms in text prompts and concepts in images, such as ‘Apoploe vesrreaitais’ meaning ‘bird’. Daras and Dimakis found that these terms often produced consistent results in isolation but also sometimes in combination with other terms in prompts. Our work is similar in that some input prompts reliably sample from the same region in latent space. We found that with prompts that include both known and unknown terms, the model often generates images from known terms, omitting unknown terms. However, when Midjourney does not ignore unknown terms, it may generate images sharing visual motifs with default images. However, unlike in a ‘secret language’, this is not a one-to-one mapping between prompt and resulting image. In our case, several unrelated prompts can produce variations of the same default image, as demonstrated in Figure 5.
@@
Because it is often unclear why the model produced a certain default image, the problem of default images touches on research in interpretability and Explainable AI (XAI). Default images reflect inherent biases and patterns encoded in the TTI model. In a sense, the default image is a nonce output and currently the only mechanism of the TTI system to let the user know that the prompt is not understood. Other generative models, such as language models, include clear, explicit ways of letting users know that something is not understood or that they cannot respond to it (e.g., refusing to answer for reasons of content policy). TTI models lack these mechanisms by default. In practice, language models internally handle such content and safety checks on Midjourney. Our prompts, however, pass these checks and obscure the fact that the model can not respond to the user prompt as expected. Instead of returning nothing, the TTI model generates default images. This could be termed a case of ostensive inferential communication (Sperber and Wilson, 1995; Wilson and Sperber, 2006), a model of communication where a speaker provides cues (ostension) to make their intended meaning clear, and the listener infers that meaning through a process of reasoning (inference). This miscommunication is also a plausible explanation as to why user satisfaction is lower when the image is different from expectations.
@@ 6.4. The Future of Default Images
We believe that in the future, the phenomenon of default images will transform for several reasons. First, the available visual material, model vocabulary, and processing power will likely continue to increase. As TTI models become more powerful, they will likely generate fewer default images in response to user prompts. Second, TTI models will continue to evolve. In particular, personalization could alter default image outputs in unpredictable ways, making it an important area for future research. However, the phenomenon of default images itself will likely not disappear entirely. We will continue to see situations where words in prompts are not understood by the generative model, where users type incoherent prompts (Kuhn et al., 2025) in error, or where users intentionally experiment with unfamiliar language, unusual phrasing, or invented terms to push the model beyond known outputs. In such cases, the model may revert to visual patterns that function as a “fallback,” even if these patterns become less recognizable as defaults in future systems.
The interplay between user prompting practices and model architecture could shape new forms of defaults. For example, as prompt engineering becomes more sophisticated, communities might discover and share emergent default patterns, treating them as creative resources rather than artifacts of model limitations. Conversely, model designers may introduce mechanisms – such as semantic interpolation or on-the-fly fine-tuning – that actively mask or diversify defaults, making them harder to detect. The prompt rewriting employed in some commercial models under-the-hood already works towards this goal.
@@
Finally, as personalization and context-awareness become standard, default images may shift from being shared phenomena to individualized experiences, tied to each user’s history, style preferences, or interaction patterns. This shift could make systematic study more challenging, but it also opens questions about how visual defaults might adapt to, or even reinforce, a user’s creative identity. In the following we review some of the limitations of our study which can inspire future work.

### v5

@@ 5.1. Default Images
Figure 4 showcases the ten default images identified with affinity diagramming in our manual study, based on the set of input prompts from Section 4.1. Some default images occurred more frequently, as depicted in Figure 6. In this figure, A1–A6 denote the different test sets, comprising 130 prompts (Section 4.1 and Appendix A). Figure 5 showcases how the same default image (in this case, ‘Lady-Birdhead’) appears in response to different prompts, with only minor visual variations. Due to Midjourney’s internal diversity sampling mechanism, the relationship between prompt and default image is many-to-many: one prompt can produce multiple different default images, and the same default image can appear in response to multiple different prompts (cf. Figure 2 and Figure 7B).
Figure 4. Set of default images. Shows the set of default images identified in our manual study, along with the descriptive labels we assign for future reference.
Figure 5. Same default image from dissimilar prompts. Demonstrates that diverse and semantically unrelated prompts still yield the same default image in Midjourney v6 and v6.1, with only minor variations.
@@
Figure 6. Default-image frequencies across test sets. Shows how often each default image appears in our six prompt groups (A1–A6), revealing consistent patterns across disparate prompt types.
@@
The default images we identified exhibit recurring motifs, reflecting the model’s aesthetic biases when faced with ambiguous or unrecognized prompts. Human figures and natural elements – such as birds, clouds, trees, and flowers – appear as primary motifs, separately or merged into strange creatures. Human-vegetal and/or human-animal fusions emerge as key subjects in default images, appearing across different atmospheres, from oneiric scenes to mythological metamorphoses. Some identified default images confirm that similarly to human dreams, “AI has an uncanny ability to combine seemingly disparate elements into sometimes sinisterly coherent representations” (PRADA).
@@
Although the literature on AI imagery attests that visual expression tends to vary according to specific TTI generators, and some note ‘default styles’ associated with different models (MANOVICH), the curated series of 10 images presented in Figure 4 reveals significant aesthetic disparities. While the five images in the lower row are reminiscent of abstract expressionist and surrealist painting, also evoking magical realism and epic fantasy films, the first five images draw from other artistic movements. As shown in figures 4 and 5, the female portrait is also an important theme, particularly prone to formal and stylistic derivations. From the solitary woman standing before a romantic landscape – in which Caspar David Friedrich’s painting resonates – to close-up representations that seem to combine the tradition of portrait painting with lyrical abstraction and contemporary fashion photography, these default images illustrate the hybrid and composite nature of AI-generated pictures. Figure 5 further emphasizes how AI-generated pictures can condense multiple artistic references into a single visual composition that, similarly to a musical piece, can then unfold in subtle variations. This series reflects the wide dissemination of individual portraits, both through the digitization and online circulation of art collections – including popular museum masterpieces – and through the massive production of born-digital pictures that fuel social media.
@@
The intriguing association of a stereotyped female figure and a bird in place of a hat in Figure 5 poses a twofold question concerning the generation of default images. Apparently, the TTI model was unable to understand the prompt or correctly recognize some of the images it was trained with. Feather hats are a recurring theme in the history of female portraiture and fashion since the Renaissance, from Titian’s Portrait of a Young Woman (c. 1536) to Rubens’ Portrait of Helena Fourment (c. 1630), from Klimt’s The Black Feathered Hat (1910) to Karl Lagerfeld’s exuberant designs for Chanel’s Spring Collection in 1992. The examples are countless, hence it seems plausible that many of these images were used to train the model. However, there seems to be a fundamental difference between human perception of visual representations of feather hats and the way computer vision ‘sees’ such images. Rather than a dataset bias, what is at stake here is a perceptual bias (OFFERT). Following a metonymical logic (i.e., taking a part to stand in for the whole), the model takes the part for the whole, assuming that the feathers correspond to the presence of a bird whose species differs across the variations of the default image. This is an insightful demonstration of “the nomadic existence to which images seem to be condemned within the historical formations of knowledge” (ALLOA).
@@ 5.2.1. Effects of seed values
We visually analyzed how different seed values affect the default images. Using different seeds with the same prompt resulted in different default images, partially sharing some of the motifs (Figure 7–A1). While there are similarities, one might not associate these images with each other. Thus, the best way to generate default images is by having a fixed seed value for all image generation runs. The images in Figure 7–A2 show default image similarities with variations in both seed values and prompts.
Figure 7. Ablation results for default-image behavior. Shows the effect of seed choice, style modifiers, known-term dominance, and model version on default images.
A: Default images remain consistent across varied seeds and prompts.
B: Style modifiers change appearance but preserve the underlying default-image motif.
C: Known concepts override unknown tokens, yielding non-default images.
D: Some Midjourney versions generate visually similar defaults from the same prompt.
@@ 5.2.2. Effects of style modifiers on default images
We found default images respond to style modifiers just like regular generated images. The same default image motifs reliably arise when applying different styles. Figure 7–B depicts how the prompt “kinkku”, which triggers default images, can be modified with different styles, like “in the style of cubism” or “oil on canvas”.
@@ 5.2.4. Effects of combining multiple default image prompts
We found that when we combined two of the default image prompts (from either the same or different sets), they would still reliably yield default images. However, as the number of default image triggering terms increased, the number of resulting default images appeared to decrease. When we combined all 20 prompts from a set, the default images identified in the main study (see Figure 4) were rare.
@@ 5.2.5. Effects of model versions
Last, we analyzed how different model versions affect the default images. In our experiments with Midjourney, both model versions 6.0 and 6.1 consistently generated very similar default images with the same prompt, as depicted in Figure 7–D. Older model versions (e.g., 5.1 and 5.2) generated default images with a more surreal or abstract aesthetic, but with similar motifs, including floating objects, human portraits intertwined with natural elements (plants or animals, or both).
Figure 8. User satisfaction with default-image outputs. Responses from our online study (N = 48) show that default images can lead to user dissatisfaction.
@@ 5.3.2. Results
Figure 8 depicts the results of our user study, showing that default images negatively affect user satisfaction, but not in all cases.
@@ 5.4. Default Images on Midjourney
Figure 9. Examples of default images and triggering prompts. A curated set of default images and their associated prompts from our dataset, illustrating the diversity of motifs and styles that recur across unrelated inputs.
@@ 5.4.1. Results
Our computational analysis of images on Midjourney resulted in the identification of 4,715 clusters of potential default images. In total, the clusters comprise 36,243 images, representing 4.84% of our dataset. This percentage refers only to our filtered dataset and parameters (i.e., less than 25 characters per prompt, lexical diversity $\geq$ 0.5, and semantic similarity $\leq$ 0.3) and should not be interpreted as representative of the overall prevalence of default images on Midjourney. The mean number of images per cluster is 7.69 (SD = 9.04, Median = 5, Min = 3, Max = 100). The lexical diversity of prompts within clusters ranges from 0.5 to 1.0 (Mean = 0.92, SD = 0.12, Median = 1.0), and the semantic similarity from 0.01 to 0.3 (Mean = 0.2, SD = 0.05, Median = 0.2). From smaller-scale experiments, we note the largest clusters were stable across multiple parameter settings. Some selected canonical examples are depicted in Figure 9 and Appendix C.
@@
  • We found evidence of short prompts, like the ones in our manual study, being in real-world use by Midjourney users. These prompts include, for instance, fantasy names, domain names, oxymorons, conceptual metaphors, abstract noun phrasing, poetic paradoxes and lyrical abstraction, surreal compound phrases, other neologisms and creative compounds, or typos. This confirms that the six categories of prompts likely triggering default images in our manual study are realistic and have practical relevance.
  • The clusters contain recurrent themes in the images, identified in our earlier manual study, such as portrait-style images, intricate cityscapes, and dream-like imagery. This includes some of the iconic Midjourney motifs, such as a person walking up an eternal staircase or road, a town on a floating island, abstract and diffuse color arrangements, and depictions of seemingly the same person.
  • Major versions of Midjourney have different sets of default images (as expected), although some versions (such as version 3 and 4) share common default image motifs. This is likely a result of training data being shared between training runs.
@@ 6.1. Default Images in Text-to-Image Generation
Default images in TTI generation are an underexplored phenomenon. We presented the first study on default images in Midjourney. When prompts are vague and contain unknown terms, Midjourney may generate default images, suggesting that the model defaults to familiar visual elements when it encounters words outside of its training data.
Our computational analysis of images on Midjourney reveals that default images are not a theoretical construct, they occur daily in response to real-world creative user prompts. Some of the images we identified in our analysis represent iconic Midjourney failure modes that many users may have encountered, especially with the early Midjourney versions. Our findings confirm that these default images are not generated randomly and, on the basis of our systematic experiments, have a relatively narrow thematic and formal range. Uncanny elements emerged in common motifs that often depicted human portraits (usually women) and animals (especially birds), along with other botanical or organic forms. Earlier versions of Midjourney have a different set of default images, often including motifs such as balloons, butterflies, fishes, floating islands, towns, or simply color pallettes and transitions (see Figure 1 and Appendix C). Although many default images appear to share a common style, this straightforward association may obfuscate a more complex reality. For instance, in model version 5, some default images appear related to surrealism and cartoonish grotesque (cf. Figure 1), with other identifiable references from art history, namely Dutch and Flemish Renaissance painting, particularly Hieronymus Bosch and Bruegel the Elder. AI-generated images are hybrid and composite, condensing multiple forms and references learned from the training dataset into a single visual output. As Wasielewski points out “style is a gradient” (Wasielewski). Therefore, an AI image is unlikely to correspond to a specific artistic movement, rather reflecting “a fluctuating set of visual characteristics with highly fuzzy boundaries” (Wasielewski).
Our findings also demonstrate that default images may negatively affect user satisfaction. In the following section, we describe the characteristics of default images in eight postulates.
@@ 6.3. Explainability and Alignment of Text-to-Image Generation Models
What does ‘absolute ownership’, ‘nallopand’, or a ‘protruding goofus’ look like? Humans could, perhaps, imagine some creative imagery for these abstract concepts and lexical inventions. However, machines struggle to produce a result (cf. Figure 1). This is not surprising, given that these terms were likely not part of the TTI models’ training data. While polysemia and ambiguity will likely continue to be an issue in the alignment of TTI models with human intent (consider, for instance, the polysemous prompt “a plant on a bank”), default images are a different phenomenon – a type of image that appears in response to prompts that are not understood by the model, revealing a misalignment between model and human intent.
@@
daras2022 experimented with how gibberish prompts are understood by the model and suspected that TTI models have developed a ‘secret language’ (daras2022). Their hypothesis was that DALL-E 2 had made ‘hidden’ associations between certain terms in text prompts and concepts in images, such as ‘Apoploe vesrreaitais’ meaning ‘bird’. daras2022 found that these terms often produced consistent results in isolation but also sometimes in combination with other terms in prompts. Our work is similar in that some input prompts reliably sample from the same region in latent space. We found that with prompts that include both known and unknown terms, the model often generates images from known terms, omitting unknown terms. However, when Midjourney does not ignore unknown terms, it may generate images sharing visual motifs with default images. However, unlike in a ‘secret language’, this is not a one-to-one mapping between prompt and resulting image. In our case, several unrelated prompts can produce variations of the same default image, as demonstrated in Figure 2.
@@
Because it is often unclear why the model produced a certain default image, the problem of default images touches on research in interpretability and Explainable AI (XAI). Default images reflect inherent biases and patterns encoded in the TTI model. In a sense, the default image is a nonce output and this output currently is the only mechanism of the TTI system to let the user know that the prompt is not understood. Other generative models, such as language models, include clear, explicit ways of letting users know that something is not understood or that they cannot respond (e.g., by refusing to answer for reasons of content policy). TTI models lack these mechanisms by default. In practice, language models internally handle such content and safety checks on Midjourney. Our prompts, however, pass these checks and obscure the fact that the model can not respond to the user prompt as expected. Instead of returning nothing, the TTI model generates default images. This could be termed a case of ostensive inferential communication (sperber1995relevance; wilson2004relevance), a model of communication where a speaker provides cues (ostension) to make their intended meaning clear, and the listener infers that meaning through a process of reasoning (inference). This miscommunication is also a plausible explanation as to why user satisfaction is lower when the image is different from expectations.
@@ 6.4. The Future of Default Images
We believe that in the future, the phenomenon of default images will transform for several reasons. First, the available visual material, model vocabulary, and processing power will likely continue to increase. As TTI models become more powerful in terms of training vocabulary, they will likely generate fewer default images in response to user prompts. Second, TTI models will continue to evolve. In particular, personalization could alter default image outputs in unpredictable ways, making it an important area for future research. As personalization and context-awareness become standard, default images may shift from being shared phenomena to individualized experiences, tied to each user’s history, style preferences, or interaction patterns. This shift could make systematic study more challenging, but it also opens questions about how visual defaults might adapt to, or even reinforce, a user’s creative identity.
However, the phenomenon of default images itself will likely not disappear entirely. We will continue to see situations where words in prompts are not understood by the generative model, where users type incoherent prompts (11084610) in error, or where users intentionally experiment with unfamiliar language, unusual phrasing, or invented terms to push the model beyond known outputs. In such cases, the model may revert to visual patterns that function as a “fallback,” even if these patterns become less recognizable as defaults in future generative systems.
Finally, the interplay between user prompting practices and model architecture could shape new forms of defaults. For example, as prompt engineering becomes more sophisticated, communities might discover and share emergent default patterns, treating them as creative resources rather than artifacts of model limitations. Conversely, model designers may introduce mechanisms – such as semantic interpolation or on-the-fly fine-tuning – that actively mask or diversify defaults, making them harder to detect. The prompt rewriting employed in some commercial models under-the-hood already works towards this goal.
6.5. Design Recommendations for Text-to-Image Systems
In light of these observations, designers of text-to-image systems can draw several lessons from default images as a signal of model limitations. First, default images point to a gap between user intent and model interpretation. Generative systems could surface this gap more transparently by clearly indicating which prompt elements are interpretable and which are not. This could take the form of hints and gentle uncertainty cues in the user interface or interactive explanations that highlight the parts of a prompt that receive little or no semantic grounding in the generative model’s internal representation. Such feedback may support user awareness without interrupting creative flow.
Second, designers could introduce mechanisms that reduce the likelihood that default images become the only response to an unfamiliar prompt. This could involve controlled fallback strategies that interpolate toward nearby known concepts while signaling that the system has done so, or the introduction of explicit “request for clarification” prompts that preserve creative freedom while maintaining communicative clarity. Some language models already ask for clarifications, and TTI generative models could be instruction-tuned to do the same. These strategies could help prevent situations where users assume intentionality in outputs that originate from unrecognized terms. On the other hand, many Midjourney users are likely not aware that some images in response to their prompts are default images. For instance, the prompt “modernika” (see Figure 10 in Appendix C) could potentially depict a fantasy figure named Modernika. Yet, this image is a default image. Whether or not the user needs to know the presence of this default image is a design choice.
Third, personalization features raise questions about how default images evolve when generative systems adapt to individual users. Designers may wish to consider whether personalized defaults should be transparent as defaults, or whether they should blend into the user’s broader creative trajectory. In other terms, designers will need to decide if users should be alerted of default images or not. Further, in creative contexts where exploration and play are central, personalized defaults could serve as material for experimentation, opening new creative pathways. Conversely, in contexts that emphasize precision, such defaults may need more explicit framing so that users can distinguish between grounded outputs and fallback patterns.
Fourth, as communities develop shared prompting practices, system designers could support ways of documenting, sharing, and surfacing these practices without reinforcing unwanted defaults. Features that let users annotate prompts, sets of prompts, and discovered quirks could guide future users and inform model improvement. Such community insight may also support internal design processes by revealing how defaults interact with the expectations and workflows of different creative user groups.
Finally, the presence of default images raises questions for future evaluation practices in TTI generation. Designers could consider incorporating default-image behavior into model diagnostics, using default images as indicators or benchmarks of blind spots in the training corpus or the prompt parser. Benchmarks that stress unknown terms or ambiguous phrasing may help identify where generative systems fail to align with user intent. Such evaluation schemes would complement traditional performance metrics and help center, but also manage, user experience in next-generation TTI models.
In the following we review some of the limitations of our study which can inspire additional future work.